OpenStudy (james1107):

There is a direct path from Gayle's house to the store. Due to construction, the direct route is closed and Gayle must travel 6 miles east and then 8 miles north in order to get to the store. How long is the direct path from Gayle's house to the store? 14 miles 12 miles 10 miles

OpenStudy (james1107):

weve got to hurry cause ive got to leave school in 10 min

OpenStudy (james1107):

and i think its the bottom one

OpenStudy (janu16):

Use the Pythagorean theorem -- a^2 + b^2 = c^2 c^2 = 6^2 + 8^2 c^2 = 36 + 64 c^2 = 100 (take the square root of 100) c = 10 miles oh youre in school? lol

OpenStudy (james1107):

yes it sucks

OpenStudy (janu16):

i m on springbreak

OpenStudy (janu16):

homw many more cause you have little time

OpenStudy (james1107):

mine is at the end of march here comes another one

OpenStudy (james1107):

@Keigh2015 say something so i can medal you

OpenStudy (keigh2015):

Okay

OpenStudy (janu16):

next one